OpenStudy (anonymous):

Find the extrema of f. f(x,y)=x-y, subject to x^2-y^2=2

OpenStudy (tkhunny):

How shall we do that? Lagrange or some other way?

OpenStudy (anonymous):

Lagrange would be best.

OpenStudy (tkhunny):

Well, set up your giant, new function: \(f(x,y,\lambda) = x - y - \lambda(x^{2} - y^{2} - 2)\). Now find the three partial derivatives. Do you need help with that?

OpenStudy (anonymous):

I'm getting (0,0,-1).

OpenStudy (tkhunny):

Waht were your partial derivatives?

OpenStudy (anonymous):

Sorry I messed up. I got \[1-\lambda 2x=0, -1-\lambda 2y=0, x^2-y^2-2=0\]

OpenStudy (tkhunny):

Unfortunately, that's not all that useful. We probably should notice that the contraint can be rewitten \(x - y = \dfrac{2}{x+y}\) except on x = -y.

OpenStudy (anonymous):

\[ f(x,y)=x-y\\ g(x,y)=x^2+y^2\\ f_x(y)=1\\ f_y(x)=-1\\ g_x(y)=2x\\ g_y(x)=2y\\ f_x=\lambda g_x\\ 2x=\frac1\lambda\\ f_y=\lambda g_y\\ -2y=\frac1\lambda\\ x=-y\\ g(x,y)=2\\ 2y^2=2\\ (1,-1,2), (-1,1,-2)\\ \]